OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = -4,096 and a4 = 64? 1 -1 4 -4

OpenStudy (anonymous):

first we need to find the ratio of the geometric series, we'll do this by writing the equation for the a4 term of a geometric series. the geometric series equation is an = a1*r^(n-1) what are the values for a1, n, and an if we just look at the given info for a4?

OpenStudy (anonymous):

a1 = -4096

OpenStudy (anonymous):

good, keep going though, if we just look at what the terms we know for when n=4 an = a1*r^(n-1) 64 = -4096*r(n-1) what would n be?

OpenStudy (anonymous):

so the 6th term is 4?

OpenStudy (anonymous):

mmm.... lets not jump ahead. look at this: an = a1*r^(n-1) if I only look at the a4 term then I know: an = 64 a1 = -4096 n = 4 take these values and plug them into an = a1*r^(n-1) what do you get?

OpenStudy (anonymous):

an = a1*r^(n-1) becomes 64 = -4096 * r^(4-1) 64 = -4096 * r^(3) the next step would be to solve for r, divide both sides by -4096

OpenStudy (anonymous):

I got -0.015 = r^(3) is that right?

OpenStudy (anonymous):

yep! now we need to get rid of the power of 3 on the r, we do this by taking the cube root of both sides, \[\sqrt[3]{-0.015} = r\]

OpenStudy (anonymous):

do u remember how to do cube roots on your calculator?

OpenStudy (anonymous):

I got -0.246 = r

OpenStudy (anonymous):

yep, just round it up to -0.25 so that is what r equals plug that into the an=a1*r(n-1) and solve for a6

OpenStudy (anonymous):

I got 4... thank you!!!!:D

OpenStudy (anonymous):

I get -1 hmm...

OpenStudy (anonymous):

a6 = -4096*(-0.25)^6

OpenStudy (anonymous):

So 64 is a4 and I multiplied 64*-0.25 = -16 -16*-0.25 = 4 and 4 would be the 6th term

OpenStudy (anonymous):

oh, i see what you did use this equations: an = a1 * r^(n-1) we found r, and know n and a1

OpenStudy (anonymous):

You got me confused, so then the answer is -1?

OpenStudy (anonymous):

geometric series equation is: an = a1 * r^(n-1) we solved for what r is -0.25 n would be 6 since we are looking for the 6th term in the sequence and a1 was given to use as -4096 and we are solving for an, which in our case would be called a6 remember that the a6 means a 'sub' 6 not a*6